OpenStudy (anonymous):

Solve for X. State any restrictions on the variables. a(bx+2)=cx-12

OpenStudy (anonymous):

Expand it out and then solve for x: abx+2a=cx-12 abx-cx=-2a-12 x(ab-c)=-2a-12 x=(-2a-12)/(ab-c)

OpenStudy (anonymous):

Restrictions: what happens if you divide by zero? Solve for ab-c = 0

OpenStudy (anonymous):

thank you